OpenStudy (anonymous):

If a player is dealt 2 cards from 2 well shuffled standard decks, what is the probability that the player will receive a blackjack?

OpenStudy (across):

Refresh my memory: How many blackjacks are there in a standard set of cards?

OpenStudy (across):

Also, how many cards are there in a standard set of cards?

OpenStudy (across):

Those two questions are essential for solving this problem.

OpenStudy (zarkon):

52 cards in a deck

OpenStudy (zarkon):

a two card black jack consists of one ace and either (K,Q,J, or 10)

OpenStudy (jamesj):

(across is being didactic)

OpenStudy (anonymous):

compute the probability that the first is an ace and the second is worth ten given that the first is an ace, the probability that the first is worth ten and the second is an ace given that the first is a ten, then add

OpenStudy (anonymous):

if you want to make your life easy do it for one suit and forget about two decks.

OpenStudy (anonymous):

so you should get \[\frac{4}{13}\times \frac{1}{12}+\frac{1}{13}\times \frac{4}{12}=\frac{2}{39}\]

OpenStudy (anonymous):

I'm going to do this in multiple posts because I just lost all my work trying to submit. There are 8 aces and 32 ten-values cards. There are two ways to get a blackjack in two draws: (i) draw an ace then draw a ten or (ii) draw a ten then draw an ace:\[\frac{8}{104}\frac{32}{103}+\frac{32}{104}\frac{8}{103}\]

OpenStudy (anonymous):

This is the same as\[=2*\frac{8}{104}\frac{32}{103} = \frac{512}{10712} \approx.0477968633\]

OpenStudy (anonymous):

or about 1 in 20.921875 times

OpenStudy (zarkon):

you can also use combinations... \[\frac{{8 \choose 1}{32\choose1}}{{104\choose2}}\]

OpenStudy (anonymous):

THANK YOU ALL SO MUCH!!